6. The Francis Company is expected to pay a dividend of D1 = $1.25 per share at the end of the year, and that dividend is expected to grow at a constant rate of 6.00% per year in the future. The company's beta is 1.70, the market risk premium is 5.50%, and the risk-free rate is 4.00%. What is the company's current stock price?
Answer #1

Required return=risk free rate+beta*market risk premium

=4+(1.7*5.5)=13.35%

Current price=D1/(Required return-Growth rate)

=1.25/(0.1335-0.06)

=$17.01(Approx).